# Calib3D: Calibrating Model Preferences for Reliable 3D Scene Understanding

### v1

@@ 3.3 DeptS: Depth-Aware Scaling for Better 3D Calibration
Observations. While previous calibration methods [86, 36, 61, 71] have shown appealing calibration performance on image-based perception tasks, their effectiveness on 3D data remains unknown. Different from RGB images, the point cloud data are unordered and texture-less, which inherits extra difficulties in feature learning [31, 29, 6]. As shown in Fig. 2, we observe a close correlation among calibration error, prediction entropy, and depth – an inherent 3D information that can be calculated based on the Cartesian coordinates $(p^{x}_{i},p^{y}_{i},p^{z}_{i})$.
@@
Depth Correlations. To further consolidate this finding, we conduct a quantitative analysis of the relation between the calibration error and depth, which is depicted in Fig. 3. We calculate the statistics of confidence and accuracy scores of all LiDAR points and then split them into 10 bins based on their depth values, where each bin corresponds to a 5-meter range. We notice from the uncalibrated result that LiDAR points with large depth values (i.e., at the middle-to-far regions of the ego-vehicle) tend to have low accuracy. However, the confidence scores of the uncalibrated model do not decrease correspondingly, which in turn causes higher calibration errors. This motivates us to design a calibration method that can resolve the over-confidence issue for LiDAR points with large depths.
@@
Depth-Aware Scaling. To fulfill the above pursuit, we propose a straightforward yet effective depth-aware scaling (DeptS) method for better calibrating 3D scene understanding models. DeptS consists of two base calibrators which are selectively used based on the prediction entropy calculated via $\hat{c}_{i}$, that is:
@@
$\hat{v}_{i}^{\text{DeptS}}=\left\{\begin{aligned} \max_{s}\sigma(\frac{\mathbf{z}_{i}}{\alpha\cdot T_{1}})^{(s)}~{},&~{}~{}~{}\text{if}-c_{i}\log(c_{i})>\eta\\ \max_{s}\sigma(\frac{\mathbf{z}_{i}}{\alpha\cdot T_{2}})^{(s)}~{},&~{}~{}~{}\text{otherwise}\end{aligned}~{},\right.$ (7)
@@
where $T_{1}$ and $T_{2}$ are the temperature parameters and satisfy $T_{1}>T_{2}$. $\eta$ is a threshold for filtering high-entropy predictions. A larger entropy value means a higher probability of being misclassified, and misclassified samples tend to encounter the over-confidence problem [13]. Therefore, we use a large $T_{1}$ to smooth the logits distribution, which will in turn reduce the confidence score.
@@
where $k_{1}$ and $k_{2}$ are two learnable parameters and satisfy $k_{1}>0$. $d_{i}$ is the depth that is calculated based on the Cartesian coordinates. In such cases, the LiDAR points with large depth values will have large $\alpha$ values, which are then used to reduce the corresponding confidence score. This in turn mitigates the over-confidence issue for points in the middle-to-far regions. The comparison between our method and the uncalibrated model in Fig. 3 shows the effectiveness of our depth-aware confidence adjustment design. As we will discuss more concretely in the following sections, DeptS contributes a stable improvement in calibrating 3D scene understanding models across a diverse spectrum of 3D datasets.
3.4 Calib3D Configurations
Aleatoric Uncertainty. 3D data are known to be highly heterogeneous due to different sensor acquisitions, placements, scene conditions, annotation protocols, pre-processing, etc. A learning-based system trained on different data sources tends to exhibit diverse predictive confidence and accuracy, especially under measurement noises. To probe such aleatoric uncertainties, a total of 10 popular 3D datasets are used in Calib3D, including 1nuScenes [29], 2SemanticKITTI [7], 3Waymo Open [99], 4SemanticPOSS [83], 5Synth4D [95], 6SemanticSTF [115], 7ScribbleKITTI [107], 8S3DIS [3], and the 9nuScenes-C and 10SemanticKITTI-C datasets from the Robo3D benchmark [55]. We hope such a comprehensive benchmark study could lay a solid foundation for the future development of reliable 3D scene understanding and beyond. Due to space limits, kindly refer to the Appendix for additional configuration details of these datasets.
Epistemic Uncertainty. The diverse spectrum of 3D scene understanding models introduces diverse factors that cope with the model uncertainty. Tailored to understand the model architecture and configuration details, Calib3D encompasses a total of 28 state-of-the-art models that have achieved promising accuracy on standard benchmarks. Based on the use of LiDAR representations, we categorize these models into five groups, i.e., 1range view [75, 21, 131, 16, 2, 54, 122], 2bird’s eye view (BEV) [129], 3voxel [18, 133, 20], 4multi-view fusion [119, 123, 101, 64, 91, 67], and 5point-based [103, 128, 112, 89, 90, 111, 121] models. Additionally, we also investigate the impact of recent 3D data augmentation techniques [113, 58, 122] and different sparse convolution backends [18, 101, 20], exhibiting design choices that matter for delivering accurate uncertainty estimates. Due to space limits, kindly refer to the Appendix for additional details.
4 Experiments
4.1 Settings
Implementation Details. We establish the Calib3D benchmark using the popular MMDetection3D [19] and OpenPCSeg [66] codebase. A total of 28 models and 10 datasets are covered in this study. We follow the default configurations when training the models, including the optimizer, learning rate, scheduler, number of training epochs, etc. The common 3D data augmentations used are random rotation, flipping, scaling, and jittering. As for the implementation of calibration methods, we follow the conventional setups from prior works. We calculate the statistics of the predictive entropy of correct/incorrect predictions and select the boundary value as the entropy threshold. We adopt a unified configuration when training the proposed DeptS and previous post-hoc calibration methods [36, 61, 110, 71]. All methods are trained for 20 epochs with a batch size of 8 and the AdamW optimizer [69]. The learning rate is set to $1$e33-3- 3 and the weight decay is $1$e66-6- 6. We use four GPUs for both training and evaluation.
Table 1: The expected calibration error (ECE, the lower the better) of state-of-the-art 3D scene understanding models on nuScenes [29] and SemanticKITTI [7]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our proposed depth-aware scaling calibrations, respectively.
@@ 4.1 Settings
Table 2: The expected calibration error (ECE) and segmentation accuracy (mIoU) of state-of-the-art 3D scene understanding models on six heterogeneous benchmarks. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibrations, respectively.
@@
Figure 4: The reliability diagrams of visualized calibration gaps from CENet [16] on the val set of SemanticKITTI [7]. UnCal, TempS, MetaC, and DeptS denote the uncalibrated, temperature, meta, and our depth-aware scaling calibrations, respectively.
Benchmark Protocols. To ensure fair comparisons, we unify the model training and evaluation configuration during benchmarking. The models are trained on the official training split of each dataset and are evaluated on data from the official val split. We reproduce the originally reported performance of the benchmarked models whenever possible, without using any extra tricks including test time augmentation, model ensembling, or fine-tuning on validation data.
Evaluation Metrics. We use the expected calibration error (ECE) metric as depicted in Eq. 2 as the main indicator in our benchmark. Following conventions, we adopt the class-wise Intersection-over-Union (IoU) and the mean Intersection-over-Union (mIoU) to measure the 3D segmentation accuracy. For robustness probing, we adopt the corruption-wise IoU scores and the mean Resilience Rate (mRR) defined in Robo3D [55] to measure the 3D robustness. Kindly refer to the Appendix for more details on these metrics.
4.2 In-Domain Uncertainty
Automotive 3D Scenes. Tab. 1 displays the calibration errors of state-of-the-art 3D scene understanding models on nuScenes [29] and SemanticKITTI [7]. We find that these models are prone to be poorly calibrated, which undermines their applicability in safety-critical contexts. We also observe similar patterns on Waymo Open [99] and SemanticPOSS [83] in Tab. 2. The calibrated uncertainties from different calibration methods [36, 61, 110, 71] exhibit promising rescues to remedy such problems. As can be seen, a simple temperature scaling [36] is useful in improving calibration. Our DeptS sets up new states for all models across all datasets, which verifies the efficacy of the proposed depth-aware scaling. A similar conclusion can be derived from Fig. 1, where DeptS holistically improves the uncertainty estimation in different regions from the LiDAR scenes.
Adverse Weather Conditions. The results of PolarNet [129], MinkUNet [18], and SPVCNN [101] on the SemanticSTF [115] dataset in Tab. 2 further showcase the importance of conducting network calibration in 3D scene understanding. Models under weather conditions fail to deliver accurate uncertainty estimates, posing challenges for real-world usage. Effective calibration using DeptS tends to provide a more reliable uncertainty estimation for safety-critical applications.
Synthetic LiDAR Data. The results on the Synth4D [95] dataset in Tab. 2 reveal the fact that models trained on synthetic data are likely to have lower calibration errors. This is because the point clouds generated by simulation engines are less complex than real-world cases. Extra attention should be drawn when adapting models trained using synthetic data to real-world applications.
Sparse Annotations. From Tab. 2, we find that models trained using weak supervisions, such as the line scribbles in ScribbleKITTI [107], tend to yield much higher calibration errors. Different from training using dense annotations, models can only learn limited knowledge from weak supervision, which inevitably leads to higher predictive uncertainties. It is thus suggested to adopt calibration methods under such cases, where the calibration errors can be effectively reduced.
Indoor 3D Scenes. The last three rows of Tab. 2 show the calibration errors on PointNet++ [90], DGCNN [111], and PAConv [121] on S3DIS [3]. As can be seen, the indoor point clouds also struggle with aleatoric and epistemic uncertainties. It becomes evident that conducting network calibration is a crucial step in achieving high levels of robustness for 3D scene understanding. Similar to other use cases in Tab. 1 and Tab. 2, our proposed DeptS serves as an effective way to reduce the gaps between confidence and predictive accuracy.
Table 3: The expected calibration error (ECE) of MinkUNet [18] under eight domain-shift scenarios from nuScenes-C and SemanticKITTI-C in the Robo3D benchmark [55]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibrations, respectively.
@@ 4.2 In-Domain Uncertainty
Reliability Diagrams. As mentioned in Sec. 3.1, the calibration gaps can be well displayed using reliability diagrams. Fig. 4 demonstrates again the superiority of DeptS; our method effectively reduces the calibration gaps (as depicted in red areas), delivering more accurate uncertainty estimates in practice than prior calibration methods [36, 71]. Additional reliability diagrams are in the Appendix.
Table 4: Ablation study on the uncertainty of 3D segmentation networks with different capacities (# of parameters).
@@
4.3 Domain-Shift Uncertainty
In addition to the in-domain scenarios, we also investigate the uncertainty estimates under the more challenging domain-shift conditions. Following the out-of-domain (OoD) configurations in Robo3D [55], the 3D scene understanding models are trained using in-domain data and are tested under OoD conditions.
Common Corruptions. The 3D data acquired from real-world scenes tend to incorporate inherent measurement noises and variants. The first four rows in Tab. 3 display such issues; the models suffer from much higher calibration errors under corruptions caused by adverse weather conditions, including fog, wet ground, and snow. The degradation from motion blur further highlights the necessity of conducting network calibration for reliable 3D scene understanding.
Sensor Failures. The beam missing, crosstalk, incomplete echo, and cross sensor scenarios in Tab. 3 reveal the vulnerability of existing 3D scene understanding models to diverse sensor failures. Compared to prior calibration methods [36, 61, 110, 71], DeptS is more stable in delivering reliable uncertainty estimates under these OoD conditions. We believe this feature is crucial towards achieving a more reliable 3D scene understanding, especially for safety-critical applications.
4.4 Ablation Study
In this section, we study several important configurations that coped closely with current 3D scene understanding research. To control variables, unless otherwise specified, we use a MinkUNet-18 [18] model with voxel size of $0.10$ m3, common augmentations, and TorchSparse [100, 102] backend on the nuScenes [29] dataset, throughout this ablation study.
Network Capacity. As has been proven in prior studies [36, 110, 71], large 2D models tend to be less calibrated than small ones. From Sec. 4.2, we observed a similar pattern for 3D scene understanding models. We find that models with fewer parameters yield lower calibration errors, albeit being less accurate. Such a phenomenon poses concerns about the development of large 3D models for safety-critical applications. Special attention should be drawn when designing models with larger capacities since they are prone to be less calibrated in practice.
3D Data Augmentations. Recent advancements in 3D data augmentations have exhibited superior 3D segmentation accuracy. In Sec. 4.2, we benchmark popular techniques, namely LaserMix [58], PolarMix [113], and FrustumMix [122], on their efficacy in uncertainty estimation. We observe large improvements in them in delivering reliable uncertain estimates, compared to their baselines.
3D Rasterization. Different rasterization resolutions tend to affect the model’s accuracy. In Sec. 4.2 and Sec. 4.2, we investigate the behaviors of models trained with different numbers of range view cells and different voxel sizes. We observe that an optimal segmentation accuracy tends to appear at a resolution that is not too high or too low. However, the calibration error poses a clear correlation with the 3D rasterization, where more range view cells or smaller voxel sizes lead to increased calibration errors and vice versa. Based on observation, it is thus suggested to take special care in configuring the training and evaluation resolutions of 3D scene understanding models with different LiDAR representations.
(a) ECE vs. mIoU
(b) SparseConv Backend
(c) LiDAR Modality
Segmentation Accuracy. We find a distinct correlation between the calibration errors and the 3D segmentation accuracy, i.e., the mIoU scores, as shown in Fig. 5(a). Similar to the observation drawn in [109], we find that a model with higher task accuracy is likely to have a relatively lower calibration error.
SparseConv Backends. We compare the behaviors of MinkUNet [18] trained using different sparse convolution backends, namely MinkowskiEngine [18], SpConv [125, 20], and TorchSparse [100, 102], and display the results in Fig. 5(b). In a general sense, SpConv [125, 20] tends to yield a higher calibration error than the other two backends. Our DeptS shows better performance across three scenarios.

### v2

@@ 3.3 DeptS: Depth-Aware Scaling for 3D Calibration
Observations. While prior calibration methods [95, 38, 66, 80] have shown appealing calibration performance on image-based perception tasks, their effectiveness on 3D data remains unknown. Unlike RGB images, point cloud data are unordered and texture-less, which inherits extra difficulties in feature learning [33, 31, 6]. As shown in Fig. 2, we observe a close correlation among calibration error, prediction entropy, and depth – an inherent 3D information derived from Cartesian coordinates $(p^{x}_{i},p^{y}_{i},p^{z}_{i})$.
@@
Depth Correlations. To consolidate this finding, we conduct a quantitative analysis of the relation between calibration error and depth (kindly refer to our Appendix). We calculate the statistics of confidence and accuracy scores of all LiDAR points and then split them into 10 bins based on their depth values, where each bin corresponds to a 5-meter range. We notice from the uncalibrated result that LiDAR points with large depth values (i.e., at the middle-to-far regions of ego-vehicles) tend to have low accuracy. However, the confidence scores of the uncalibrated model do not decrease correspondingly, leading to higher calibration errors. This motivates us to design a method that can resolve the over-confidence issue for LiDAR points with large depths.
@@
Depth-Aware Scaling. To fulfill the above pursuit, we propose a simple yet effective depth-aware scaling (DeptS) method for better calibrating 3D scene understanding models. DeptS employs two base calibrators which are selectively used based on prediction entropy calculated using $\hat{c}_{i}$, which is formulated as follows:
@@
$\hat{v}_{i}^{\text{DeptS}}=\begin{cases}\max_{s}\sigma(\frac{\mathbf{z}_{i}}{\alpha\cdot T_{1}})^{(s)}~{},&\text{if }-c_{i}\log(c_{i})>\eta\\ \max_{s}\sigma(\frac{\mathbf{z}_{i}}{\alpha\cdot T_{2}})^{(s)}~{},&\text{otherwise}\\ \end{cases}~{},$ (7)
@@
where $T_{1}$ and $T_{2}$ are temperature parameters and satisfy $T_{1}>T_{2}$. The threshold $\eta$ filters high-entropy predictions. Higher entropy indicates a greater likelihood of misclassification, which is often associated with over-confidence [13]. Therefore, we use a larger $T_{1}$ to smooth the logits distribution, which will in turn reduce the confidence score.
@@
where $k_{1}$ and $k_{2}$ are learnable parameters with $k_{1}>0$. $d_{i}$ is the depth that is calculated based on the Cartesian coordinates. In this way, LiDAR points with large depth values will have large $\alpha$ values, which are then used to reduce the corresponding confidence score. This in turn mitigates the over-confidence issue for points in the middle-to-far regions. The comparison between our method and the uncalibrated model exhibits the effectiveness of our depth-aware confidence adjustment design. As we will discuss more concretely in the following sections, DeptS contributes a stable improvement in calibrating 3D scene understanding models across a diverse spectrum of 3D datasets.
Table 1: The expected calibration error (ECE, the lower the better) of state-of-the-art 3D scene understanding models on the validation sets of the nuScenes [31] and SemanticKITTI [7] datasets. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our proposed depth-aware scaling calibration methods, respectively.
@@
3.4 Benchmark Configurations
Aleatoric Uncertainty. 3D data are inherently diverse due to variations in sensor types, placements, and scene conditions. A learning-based system trained on such heterogeneous data often exhibits differing levels of confidence and accuracy, especially under measurement noise. To explore aleatoric uncertainty, Calib3D includes 10 popular 3D datasets: 1nuScenes [31], 2SemanticKITTI [7], 3Waymo Open [108], 4SemanticPOSS [92], 5Synth4D [104], 6SemanticSTF [125], 7ScribbleKITTI [116], 8S3DIS [3], and 9nuScenes-C and 10SemanticKITTI-C from Robo3D [59]. This comprehensive study aims to provide a foundation for developing reliable 3D scene understanding models. For additional dataset details, please refer to the Appendix.
Epistemic Uncertainty. The diverse range of 3D models introduces factors that influence model uncertainty. Calib3D includes 28 state-of-the-art models with promising performance on standard benchmarks. Based on LiDAR representations, these models are categorized into five groups: 1range view [84, 23, 142, 18, 2, 58, 133], 2bird’s eye view (BEV) [141], 3voxel [20, 144, 22], 4multi-view fusion [131, 134, 110, 71, 100, 75], and 5point-based models [112, 139, 122, 98, 99, 120, 132]. We also examine the impact of 3D data augmentation techniques [123, 62, 133] and sparse convolution backends [20, 110, 22], identifying key design factors for accurate uncertainty estimates. For additional details, please refer to the Appendix.
4 Experiments
4.1 Settings
Implementation Details. The Calib3D benchmark is built using the popular MMDetection3D [21] and OpenPCSeg [74] codebases, covering a total of 28 models and 10 datasets. We adhere to default configurations for training the models, including the optimizer, learning rate, scheduler, number of training epochs, etc. Common 3D data augmentations, such as random rotation, flipping, scaling, and jittering, are also applied. For the calibration methods, we follow the conventional setups from prior works. We calculate the predictive entropy statistics for correct/incorrect predictions to select the boundary value as the entropy threshold. Both the proposed DeptS and previous post-hoc calibration methods [38, 66, 119, 80] are trained under unified configurations. All methods are trained for 20 epochs with a batch size of 8, using the AdamW optimizer [78]. The learning rate is set to $1$e33-3- 3, and the weight decay is $1$e66-6- 6. We use four GPUs for both training and evaluation.
Benchmark Protocols. To ensure fair comparisons, we unify model training and evaluation configurations during benchmarking. Models are trained on the official training split of each dataset and evaluated on the val split. We reproduce the originally reported performance without using any extra tricks including test time augmentation, model ensembling, or fine-tuning on validation data.
Evaluation Metrics. The expected calibration error (ECE) metric, as depicted in Eq. 2, is the primary benchmark indicator. We also use class-wise Intersection-over-Union (IoU) and mean IoU (mIoU) to measure 3D segmentation accuracy. For robustness probing, we adopt corruption-wise IoU scores and the mean Resilience Rate (mRR) from Robo3D [59] to measure the 3D robustness. Kindly refer to the Appendix for more details on these metrics.
4.2 In-Domain Uncertainty
Automotive 3D Scenes. Tab. 1 shows the calibration errors of state-of-the-art 3D scene understanding models on nuScenes [31] and SemanticKITTI [7]. We observe that these models are often poorly calibrated, raising concerns about their reliability in safety-critical contexts. Similar patterns are seen on Waymo Open [108] and SemanticPOSS [92] in Tab. 2. Different calibration methods [38, 66, 119, 80] show promising results in addressing these issues, with temperature scaling [38] being particularly effective. Our DeptS sets new benchmarks across all models and datasets, demonstrating the benefits of depth-aware scaling. As highlighted in Fig. 1, DeptS holistically improves uncertainty estimation in various regions of LiDAR scenes.
Adverse Weather Conditions. The results of PolarNet [141], MinkUNet [20], and SPVCNN [110] on the SemanticSTF [125] dataset, shown in Tab. 2, underscore the importance of network calibration in 3D scene understanding. Weather conditions pose challenges to accurate uncertainty estimates, complicating real-world applications. Effective calibration with DeptS provides more reliable uncertainty estimations, crucial for safety-critical use cases.
Synthetic LiDAR Data. The results on the Synth4D [104] dataset in Tab. 2 suggest that models trained on synthetic data tend to have lower calibration errors, likely due to the less complex nature of simulated point clouds compared to real-world cases. Extra caution is advised when transferring these models to real-world applications.
Sparse Annotations. Tab. 2 also shows that models trained with weak supervision, such as the line scribbles in ScribbleKITTI [116], tend to exhibit higher calibration errors. Compared to dense annotations, weak supervision restricts the model’s learning capacity, leading to increased predictive uncertainties. It is thus suggested to adopt calibration methods under such cases to effectively reduce these errors.
Indoor 3D Scenes. The last three rows of Tab. 2 show calibration errors for PointNet++ [99], DGCNN [120], and PAConv [132] on the S3DIS [3] dataset. Indoor point clouds also suffer from aleatoric and epistemic uncertainties, emphasizing the importance of network calibration for robust 3D scene understanding. As seen in Tab. 1 and Tab. 2, DeptS effectively narrows the gap between confidence and predictive accuracy in these challenging environments.
Reliability Diagrams. As discussed in Sec. 3.1, calibration gaps are well illustrated through reliability diagrams. Fig. 3 highlights the effectiveness of DeptS in reducing these gaps (depicted in red areas), delivering more accurate uncertainty estimates in practice than prior calibration methods [38, 80]. Additional reliability diagrams are in the Appendix.
Figure 3: The reliability diagrams of visualized calibration gaps from CENet [18] on SemanticKITTI [7]. UnCal, TempS, LogiS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, meta, and our depth-aware scaling calibration methods, respectively.
@@ 4.2 In-Domain Uncertainty
Table 2: The expected calibration error (ECE, the lower the better) and segmentation accuracy (mIoU, the higher the better) of state-of-the-art 3D scene understanding models on the validation sets of six heterogeneous benchmarks. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibration methods, respectively.
@@
Table 3: The expected calibration error (ECE the lower the better) of the MinkUNet [20] model under eight domain-shift scenarios from the nuScenes-C and SemanticKITTI-C datasets in the Robo3D benchmark [59]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibration methods, respectively.
@@
4.3 Domain-Shift Uncertainty
Beyond in-domain scenarios, we also explore uncertainty estimates under more challenging domain-shift conditions. Following the out-of-domain (OoD) settings from Robo3D [59], we train 3D scene understanding models on in-domain data and test them under OoD conditions.
Common Corruptions. Real-world 3D data often include inherent measurement noise and variations. The first four rows of Tab. 3 illustrate these issues, showing that models experience significantly higher calibration errors under corruptions caused by adverse weather conditions, including fog, wet ground, and snow. The degradation from motion blur further emphasizes the importance of network calibration for reliable 3D scene understanding.
(a) ECE vs. mIoU
(b) Backend
(c) LiDAR Modality
Sensor Failures. The beam missing, crosstalk, incomplete echo, and cross sensor scenarios in Tab. 3 expose the vulnerability of existing 3D scene understanding models to various sensor failures. Compared to prior calibration methods [38, 66, 119, 80], DeptS is more stable in providing uncertainty estimates under these OoD conditions. This robustness is essential for achieving more reliable 3D scene understanding, particularly in safety-critical applications.
4.4 Ablation Study
In this section, we study several key settings that coped closely with current 3D scene understanding research. To control variables, unless otherwise specified, we use a MinkUNet-18 [20] model with voxel size of $0.10$ m3, common augmentations, and TorchSparse [109, 111] backend on the nuScenes [31] dataset throughout this ablation study.
Table 4: Ablation study on the uncertainty of 3D segmentation networks with different model capacities (# of parameters).
@@ 4.4 Ablation Study
Network Capacity. Prior studies [38, 119, 80] have shown that larger 2D models tend to be less calibrated than smaller ones. From Tab. 4, we observed a similar trend in 3D scene understanding models. Models with fewer parameters exhibit lower calibration errors, albeit being less accurate. This raises concerns about the development of large 3D models for safety-critical applications. Special attention should be drawn when designing models with larger capacities since they are prone to be less calibrated in practice.
3D Data Augmentations. Recent advancements in 3D data augmentations have exhibited superior 3D segmentation accuracy. In Tab. 5, we benchmark popular techniques, namely LaserMix [62], PolarMix [123], and FrustumMix [133], on their efficacy in uncertainty estimation. We observe large improvements in them in delivering reliable uncertain estimates, compared to their baselines.
@@
3D Rasterization. The resolution of 3D rasterization impacts both accuracy and calibration error. As seen in Tab. 6 and Tab. 7, Optimal segmentation accuracy typically occurs at moderate resolutions. However, calibration error poses a clear correlation with the 3D rasterization, where more range view cells or smaller voxel sizes lead to increased calibration errors and vice versa. Careful consideration is required when configuring resolutions for training and evaluation across different LiDAR representations.
Segmentation Accuracy. We find a distinct correlation between calibration errors and 3D segmentation accuracy, i.e., mIoU scores, as shown in Fig. 4(a). Similar to the observation drawn in [118], we find that a model with higher task accuracy is likely to have a relatively lower calibration error.
SparseConv Backends. We compare the behaviors of MinkUNet [20] trained using different sparse convolution backends, i.e., MinkowskiEngine [20], SpConv [136, 22], and TorchSparse [109, 111], and display the results in Fig. 4(b). In a general sense, SpConv [136, 22] tends to yield a higher calibration error than the other two backends. Our DeptS shows better performance across three scenarios.